For what value(s) of k, is the function f defined by
f(x) = {sin (1 + x) - sin(1- x) x not equal to 0
{k cos 1, x= 0
Continuous at x = 0 ?
1
Expert's answer
2018-11-19T08:46:52-0500

The given function

"f(x)=\begin{cases}\sin(1+x)-\sin(1-x)&x\neq0\\k\cos1&x=0\end{cases}"

is continuous, if

"\lim_{x \to 0}(\sin(1+x)-\sin(1-x))=\lim_{x \to 0}k\cos1"

Let us solve this equation with respect to k

"\sin1-\sin1=k\cos1""k\cos1=0""k=0"

Thus, the function f(x) is continuous for k = 0.